OpenStudy (anonymous):

Find the solution to the following equation by transforming it into a perfect square trinomial. x2 – 10x = 39

OpenStudy (anonymous):

add the square of the half of 10 to both sides

OpenStudy (mathmale):

Starting with x2 – 10x = 39, rewrite it (in Equation Editor or in the Draw utility) as\[1x^2 – 10x = 39\]Compare the left side of this result to the general form of the quadratic polynomial: \[y=ax^2 +bx+c\]

OpenStudy (mathmale):

Comparing, you'll see that a = ?? and b = ??

OpenStudy (mathmale):

Example: The given equation is \[1x^2-10x=39\]and the first two terms of the general quadratic poly. are \[ax^2+bx\] Comparing \[1x^2~\to ~ax^2\] you can easily see that a = 1. Your turn: please determine the value of b.